OpenStudy (anonymous):

Find the difference in the volume and total area of a cylinder with both a radius and height of 1. r = 1, h = 1 The number of sq units of the total area exceeds the number of cu. units in the volume by

OpenStudy (aaronq):

Find the difference means subtract one from the other (the smaller from the larger since a negative volume or surface area doesn't make sense) V=pi*r^2*h SA=2*pi*r^2+2*pi*r*h